University of Niš Faculty of Science and Mathematics

5th International Symposium on

MATHEMATICAL ANALYSIS AND ITS APPLICATIONS

MAA5

NIŠKA BANJA, October, 2-6, 2002

First announcement

Following the tradition of the previous four symposia on Mathematical Analysis and its Applications, under the auspices of the Union of the Mathematical Societies of Yugoslavia, the Faculty of Science and Mathematics, University of Niš, organizes the Fifth Symposium in the spa of Niška Banja, 10 km from Niš, from Wednesday 2nd October to Sunday 6th October, 2002. The scientific program consists of lectures (45 minutes) and research talks (20 minutes). The program of the Symposium and the book of abstracts will be available at the Symposium. Submitted papers will be published in the Proceedings of the Symposium after the usual refereeing process.

Sections:

1. Functional Analysis and Operator Theory

2. Real and Complex Analysis

3. Differential Equations

4. Applications of Analysis

Pogram committee  Conference information

and contacts

V. Rakočević, Niš, Chair

Z. Kadelburg,Beograd  MAA5 Organizing committee
Ž. Pavićević, Podgorica  Faculty of Science and Mathematics
S. Pilipović, Novi Sad  University of Niš

Višegradska 33

Organizing committee  18 000 Niš, Serbia, Yugoslavia
D. Djordjević, Chair  Fax: (+381 18) 533014
D. Cvetković  Phone: (+381 18) 324 492
S. Janković  E-mail: maa5@pmf.pmf.ni.ac.yu
I. Jovanović  www.pmf.ni.ac.yu

M. Jovanović

E. Malkowsky,  Important dates:

J. Manojlović

M. Milosavljević  Deadline for the submission of abstracts:
S. Tričković  May 15, 2002
S. Vujović  Preliminary registration: July 15, 2002
MAA5 October 2-6, 2002.

Registration form MAA5

Name/Names Institution Address E-mail Phone Fax Section: 1 2 3 4 Hotel accommodation: Yes / No The title of the talk

Abstract